OpenStudy (tiffany_rhodes):

How to find the antiderivative of:

OpenStudy (tiffany_rhodes):

\[\int\limits_{1}^{x}\sqrt{1+(t^2-(1/(4t^2))^2} dt\]

OpenStudy (tiffany_rhodes):

I'm sorry it's (t^2 -(1/(4t^2)))^2

zepdrix (zepdrix):

\[\large \int\limits_1^x \sqrt{1+\left(t^2-\frac{1}{4t^2}\right)^2}\] It's a little hard to read the way you've written it out. Is this correct?

zepdrix (zepdrix):

Woops I forgot my dt :D whatev.

OpenStudy (tiffany_rhodes):

lol yes! Sorry for the confusion :)

zepdrix (zepdrix):

Mmk so let's work from the inside out.\[\large t^2-\frac{1}{4t^2}\qquad = \qquad \frac{4t^4-1}{4t^2}\]I got a common denominator and combined them.

zepdrix (zepdrix):

Now we have to square this..? Hmm

zepdrix (zepdrix):

Squaring that term gives us this, \[\large \frac{16t^8-8t^4+1}{16t^4}\] Any confusion on how I got that?

OpenStudy (tiffany_rhodes):

nope, I got it!

zepdrix (zepdrix):

So I guess that leaves us here, \[\large \int\limits\limits_1^x \sqrt{1+\frac{16t^8-8t^4+1}{16t^4}}\quad dt\] Once again, we'll want to get a common denominator between that 1 and the fraction.

zepdrix (zepdrix):

\[\large \int\limits\limits\limits_1^x \sqrt{\frac{16t^4}{16t^4}+\frac{16t^8-8t^4+1}{16t^4}}\quad dt\]

zepdrix (zepdrix):

\[\large \int\limits\limits\limits\limits_1^x \sqrt{\frac{16t^8+8t^4+1}{16t^4}}\quad dt\]

OpenStudy (tiffany_rhodes):

hmmm okay. Now I just need to find the antidertivative of that?

zepdrix (zepdrix):

I have a feeling what we're going to have to do is probably... complete the square in the numerator, and then apply a trig sub. Have you done work with trigonometric substitutions?

zepdrix (zepdrix):

We can take the root of the denominator really nicely. Let's do that, giving us,\[\large \int\limits_1^x \frac{1}{4t^2}\sqrt{16t^8+8t^4+1}\quad dt\]

OpenStudy (tiffany_rhodes):

No I haven't. I think we are skipping that section.

zepdrix (zepdrix):

Ah ok, maybe that's not the next step then. Hmm what to do next. Thinkinggg

OpenStudy (zarkon):

factor

OpenStudy (zarkon):

under the radical

OpenStudy (zarkon):

\[(4t^4+1)^2\]

zepdrix (zepdrix):

Oh it's already a perfect square under the radical? Blah, I dunno why I didn't notice that lol.

zepdrix (zepdrix):

\[\large \int\limits\limits_1^x \frac{1}{4t^2}\sqrt{(4t^2+1)^2}\quad dt\]Hmm that works out nicely :)

zepdrix (zepdrix):

Do you think you can solve it from here tiffany? :D

OpenStudy (tiffany_rhodes):

yes, thank you both! :)